OpenStudy (anonymous):

The perimeter of a rectangle is 34 feet. Find the width of the rectangle if the length is 5 feet more than twice the width.

OpenStudy (anonymous):

let , width=w therefore, length(l)=2w+5 now, perimeter =2(l+b) 34=2(l+b) 34=2(2w+5+w) 34=2(3w+5) 34/2=3w+5 17=3w+5 17-5=3w 12=3w 12/3=w 4=w so, width =4 feet

OpenStudy (anonymous):

Thank you, I just wanted to check my answer and it was correct :-)

OpenStudy (anonymous):

:)